# Minibatch Optimal Transport and Perplexity Bound Estimation in Discrete Flow Matching

### v3

@@ Appendix C Additional Experimental Results
The foundational architecture of the model we use is based on the diffusion transformer paradigm outlined by Peebles & Xie (2023), which adapts the classic encoder-only transformer structure, such as that introduced in Vaswani et al. (2017); Devlin et al. (2019), by incorporating time-based conditioning. This approach introduces slight architectural modifications, notably the use of rotary positional embeddings as described in Su et al. (2024). Due to the addition of time conditioning, the model’s parameter count is approximately 5% higher than that of a typical transformer (e.g., GPT-2). Tokenization and dataset splits are kept consistent with previous work to maintain comparability and minimize confounding variables.
The architecture comprises 12 transformer layers, each equipped with 12 attention heads and a hidden dimensionality of 768, matching the configuration commonly referred to as GPT-2. A dedicated conditioning dimension of 128 is used to capture temporal features essential to the diffusion process. It utilizes conventional scaled dot-product attention and applies a dropout rate of 0.1 to counter overfitting.
Regarding the training setup for OWT experiments, each model was trained with sequence lengths of 128 using a single H200 GPU. The vocabulary includes 50,257 tokens, and the training batch size is fixed at 512. The training schedule encompasses 400,000 steps, and takes 44 hours in the standards case, which increases to 45 when using OT.
The OpenWebText dataset serves as the primary training corpus with local data storage employed to reduce latency. In all cases, we use the schedule $k_{t}=\epsilon+(1-\epsilon)t$ with parameter $\epsilon=0.001$, consistent with settings from Lou et al. (2024). Evaluation samples are generated using 128 or 1024 steps.
Optimization is handled via the AdamW algorithm, set with a learning rate of 3e-4, beta values of (0.9, 0.999), and an epsilon of 1e-8. No weight decay is used, favoring pure learning rate dynamics. A warm-up phase of 2,500 steps is included to enhance training stability, and gradient clipping is applied at a value of 1.
@@ C.2 Training Bound Comparisons
c) DFM-N is the same as DFM-S but multiplied by (1t)(1-t): $\int_{0}^{1}\sum_{x_{1},x_{0}}\pi(x_{1},x_{0})\sum_{x_{t}}p_{t|1,0}(x_{t}|x_{1},x_{0})\sum_{i=1}^{L}-\delta_{m}(x_{t}^{i})\log{p_{1|t}^{i}(x^{i}_{1}|x_{t};\theta)}dt.$
The model architecture in all cases is identical in design as the one in Section 5.1, but here we use the GPT2 tokenizer and to match related work, we train on OWT (Gokaslan & Cohen, 2019) for 400K steps with batch size of 512, sequence length of 128. For ‘DFM-S’, our bound becomes the MD4 of Shi et al. (2024) (see Appendix A.4). The bound is tested on the test sets found in Lou et al. (2024), more precisely: 1BW, LAMBADA, PTB, Wikitext2 and Wikitext103 (Chelba et al., 2013; Paperno et al., 2016; Marcus et al., 1993; Merity et al., 2016). In addition, we compare against SEDD of Lou et al. (2024). The results can be below in Table 6.
@@
In Table 7 and 8, we provide the perplexity bound results on the five test sets for the models described in Section 5.3.
@@
DFM-B $\bm{184.81}$ 211.66 723.15 207.73 230.87
DFM-B-OT 190.21 $\bm{204.16}$ $\bm{654.88}$ $\bm{204.22}$ $\bm{222.42}$
@@
Table 8: DFM‑MMLM perplexity bound results comparing normal training vs OT. Sinkhorn Solver at test time.
@@
DFM‑MMLM 68.65 $\bm{68.38}$ 204.17 $\bm{68.68}$ 85.09
DFM‑MMLM-OT $\bm{68.63}$ 69.33 $\bm{204.06}$ 69.21 $\bm{83.45}$
Table 9: DFM‑MMLM perplexity bound results comparing normal training vs OT. Exact OT solver at test time.
Dataset Lambada Wiki2 PTB Wiki3 LM1B
DFM-O 71.90 71.23 221.62 70.80 82.60
DFM‑MMLM 68.65 $\bm{68.38}$ 204.17 $\bm{68.68}$ 85.09
DFM‑MMLM-OT $\bm{68.37}$ 69.10 $\bm{202.63}$ 69.02 $\bm{83.05}$
@@
In what follows we present the full generative perplexity results of the experiments described in Section 5.3. That is, we show show resutls when Llama is used as a judge, the entropy values and standard deviations. Such resulrs can be found in tables 10, 11 and 12.
@@
C.5 Tightness of bounds
@@
which shows that it depends on the ratios of induced pathwise probabilities under the model, which are intractable. Unfortunately, this makes this term difficult ot estimate in practice, as computing these ratios would require summing over all possible trajectories that reach a given state at time $t$, which is infeasible due to the uncountably infinite number of such paths.
@@ C.7 Correlation Between the Perplexity Bound and Generative Perplexity Results
For each model, we compute two metrics: (1) the average perplexity bound across 5 test sets, and (2) the generative perplexity measured by GPT2-Large with 1024 generation steps. The Pearson correlation coefficient between these metrics is 0.96, indicating very strong correlation. If we normalize the columns of the perplexity results in order to equalize the contribution of each testing set, then the Pearson correlation coefficient changes to 0.958.
@@ C.8 Ablation Experiments
Ablation 1: OT Solver. We replace the Sinkhorn solver with exact OT for multimasked flows. Results appear in Tables 14 and 15. We did not notice differences in training time.
Table 14: Exact OT (MMLM) — 512 OT, 512 dff
@@
Ablation 2: OT Batch Size. We increase the OT batch size from 512 to 4096 while maintaining the diffusion batch size at 512, performing 8 parameter updates per OT batch. We continue using exact OT. Results appear in Tables 16 and 17.
Table 16: Generative Perplexity: Exact OT (MMLM) — 4096 OT, 512 dff
@@ D.2 Continuous-Time Markov Chains Over Finite-State Spaces (Discrete Diffusion)
For the choice $Q_{t}=\sigma^{{}^{\prime}}(t)Q$, where $\sigma$ is monotonically increasing, $\sigma(0)=0$ and $\lim_{t\rightarrow 1}\sigma(t)=T$, we have $s_{t}=e^{\sigma(t)Q}s_{0}$. Matrices $Q$ must satisfy the properties of transition-rate matrices (Suhov & Kelbert, 2008), that is, they have non-negative non-diagonal entries, and the elements in each column add to 0. The choice for $Q$ is made such that: $\bm{s}_{1}$ is an easy reference distribution to sample from and the matrix exponential $e^{\sigma(t)Q}$ is easy to calculate (Austin et al., 2021; Campbell et al., 2022). Unfortunately, these conditions greatly restrict the design space in this framework.
@@ 31 James Wood, 1925-2002
Listing 6: Generated text from the DFM-MMLM model with multi-mask source distribution and normal training, with sequence length L=128.
@@
Listing 7: Generated text from the DFM-MMLM-OT model with multi-mask source distribution and minibatch-OT training, with sequence length L=128.

### v4

@@ Appendix C Additional Experimental Results
The foundational architecture of the model we use is that of Lou et al. (2024); Haxholli et al. (2025) which is based on the diffusion transformer paradigm outlined by Peebles and Xie (2023), which adapts the classic encoder-only transformer structure, such as that introduced in Vaswani et al. (2017); Devlin et al. (2019), by incorporating time-based conditioning. This approach introduces slight architectural modifications, notably the use of rotary positional embeddings as described in Su et al. (2024). Due to the addition of time conditioning, the model’s parameter count is approximately 5% higher than that of a typical transformer (e.g., GPT-2). Tokenization and dataset splits are kept consistent with previous work to maintain comparability and minimize confounding variables.
The architecture comprises 12 transformer layers, each equipped with 12 attention heads and a hidden dimensionality of 768, matching the configuration commonly referred to as GPT-2. A dedicated conditioning dimension of 128 is used to capture temporal features essential to the flow process. It utilizes conventional scaled dot-product attention and applies a dropout rate of 0.1 to counter overfitting.
Regarding the training setup for OWT experiments, each model was trained with sequence lengths of 128 using a single GH200 GPU. The vocabulary includes 50,257 tokens, and the training batch size is fixed at 512. The training schedule encompasses 400,000 steps, and takes 44 hours in the standard case, which increases to 45 when using OT.
The OpenWebText dataset serves as the primary training corpus with local data storage employed to reduce latency. Optimization is handled via the AdamW algorithm, set with a learning rate of 3e-4, beta values of (0.9, 0.999), and an epsilon of 1e-8. No weight decay is used, favoring pure learning rate dynamics. A warm-up phase of 2,500 steps is included to enhance training stability, and gradient clipping is applied at a value of 1.
In the OT-EMA experiments we employ an Exponential Moving Average (EMA) of the model parameters with a high target decay rate of $\beta=0.99999$. To prevent initialization bias, the decay is dynamically ramped up at step $t$ according to $\min(\beta,\frac{1+t}{10+t})$. Such EMA parameters are only used when calculation minibatch OT between embeddings.
@@ C.2 Training Bound Comparisons
c) DFM-N is the same as DFM-S but multiplied by (1t)(1-t):
$\int_{0}^{1}\sum_{x_{1},x_{0}}\pi(x_{1},x_{0})\sum_{x_{t}}p_{t|1,0}(x_{t}|x_{1},x_{0})\sum_{i=1}^{L}-\delta_{m}(x_{t}^{i})\log{p_{1|t}^{i}(x^{i}_{1}|x_{t};\theta)}dt.$
The model architecture in all cases is identical in design as the one in Section 5.1, but here we use the GPT2 tokenizer and to match related work, we train on OWT (Gokaslan and Cohen, 2019) for 400K steps with batch size of 512, sequence length of 128. For ‘DFM-S’, our bound becomes the MD4 of Shi et al. (2024) (see Appendix A.4). The bound is tested on the test sets found in Lou et al. (2024), more precisely: 1BW, LAMBADA, PTB, Wikitext2 and Wikitext103 (Chelba et al., 2013; Paperno et al., 2016; Marcus et al., 1993; Merity et al., 2016). In addition, we compare against SEDD of Lou et al. (2024). The results can be below in Table 6.
@@
In Table 7 and 8, we provide the perplexity bound results on the five test sets for the models trained with cross entropy described in Section 5.3.
@@
DFM-B 184.81 211.66 723.15 207.73 230.87
DFM-B-OT 190.21 204.16 654.88 204.22 222.42
DFM-B-Exact 168.02 189.16 676.29 191.16 209.75
DFM-B-Exact-EMA $\boldsymbol{167.10}$ $\boldsymbol{175.09}$ $\boldsymbol{618.78}$ $\boldsymbol{176.53}$ $\boldsymbol{204.14}$
@@
Table 8: DFM‑MMF perplexity bound results comparing normal training vs OT. Sinkhorn Solver at test time.
Dataset Lambada Wiki2 PTB Wiki3 LM1B
DFM-O 71.90 71.23 221.62 70.80 $\boldsymbol{82.60}$
DFM‑MMF 68.65 $\boldsymbol{68.38}$ 204.17 $\boldsymbol{68.68}$ 85.09
DFM‑MMF-OT $\boldsymbol{68.63}$ 69.33 $\boldsymbol{204.06}$ 69.21 83.45
Table 9: DFM‑MMF perplexity bound results comparing normal training vs OT. Exact OT solver at test time.
@@
DFM‑MMF 68.65 68.38 204.17 68.68 85.09
DFM‑MMF-OT 68.37 69.10 202.63 69.02 83.05
DFM‑MMF-Exact $\boldsymbol{66.27}$ 67.60 $\boldsymbol{197.05}$ 67.62 82.98
DFM‑MMF-Exact-EMA 68.62 $\boldsymbol{67.37}$ 213.14 $\boldsymbol{67.32}$ $\boldsymbol{82.34}$
@@
In what follows we present the full generative perplexity results of the experiments described in Section 5.3. That is, we show show resutls when Llama is used as a judge, the entropy values and standard deviations (calculated over generated samples, not over runs). Such resulrs can be found in tables 10, 11 and 12.
@@
DFM-B-Exact 335.14 235.15 206.11 194.23 188.85 180.91
Standard deviation $\pm 2.03$ $\pm 1.42$ $\pm 1.31$ $\pm 1.38$ $\pm 1.22$ $\pm 1.28$
DFM-B-Exact-EMA 302.76 208.42 182.40 169.89 164.42 159.87
Standard deviation $\pm 1.85$ $\pm 1.34$ $\pm 1.19$ $\pm 1.03$ $\pm 0.99$ $\pm 1.04$
@@
DFM-MMF-Exact 518.86 281.39 199.68 168.12 153.51 141.46
Standard Deviation $\pm$2.99 $\pm$1.57 $\pm$1.19 $\pm$0.97 $\pm$0.95 $\pm$0.98
DFM-MMF-Exact-EMA $480.61$ $259.44$ $187.68$ $156.52$ $143.08$ $132.55$
Standard deviation $\pm 2.66$ $\pm 1.38$ $\pm 1.10$ $\pm 0.90$ $\pm 0.79$ $\pm 0.82$
@@
DFM-B-Exact 387.05 276.84 248.27 232.19 227.61 218.93
Standard deviation $\pm 2.56$ $\pm 1.88$ $\pm 1.90$ $\pm 1.70$ $\pm 1.75$ $\pm 1.86$
DFM-B-Exact-EMA 351.51 246.36 220.30 204.70 197.86 193.70
Standard deviation $\pm 2.24$ $\pm 1.68$ $\pm 1.59$ $\pm 1.48$ $\pm 1.44$ $\pm 1.50$
@@
DFM-MMF-Exact 602.34 337.54 245.29 207.90 190.66 177.82
Standard Deviation $\pm$3.47 $\pm$2.03 $\pm$1.64 $\pm$1.39 $\pm$1.35 $\pm$1.41
DFM-MMF-Exact-EMA 555.42 311.88 230.42 194.47 178.12 164.78
Standard Deviation $\pm$3.02 $\pm$1.81 $\pm$1.5 $\pm$1.26 $\pm$1.25 $\pm$1.23
@@
DFM-MMF-Exact 6.35 6.31 6.28 6.26 6.25 6.24
DFM-MMF-Exact-EMA 6.35 6.31 6.28 6.26 6.25 6.23
Standard Deviation $\pm$0.001 $\pm$0.001 $\pm$0.001 $\pm$0.001 $\pm$0.002 $\pm$0.002
C.5 Tightness of Bounds Evaluation
@@
which shows that it depends on the ratios of induced pathwise probabilities under the model, which are intractable. Unfortunately, this makes this term difficult to estimate in practice, as computing these ratios would require summing over all possible trajectories that reach a given state at time $t$, which is infeasible due to the uncountable infinite number of such paths.
@@ C.7 Correlation Between the Perplexity Bound and Generative Perplexity Results
For each model in Table 3, we compute two metrics: (1) the average perplexity bound across 5 test sets, and (2) the generative perplexity measured by GPT2-Large with 1024 generation steps. The Pearson correlation coefficient between these metrics is 0.933, indicating very strong correlation. If we normalize the columns of the perplexity results in order to equalize the contribution of each testing set, then the Pearson correlation coefficient changes to 0.932.
@@ C.8 Ablation Experiments
Ablation 1: OT Solver. We replace the Sinkhorn solver with exact OT for multimasked flows (MMF). Results appear in Tables 14 and 15. We did not notice differences in training time.
Table 14: Exact OT (MMF) — 512 OT, 512 dff
@@
Ablation 2: OT Batch Size. We increase the OT batch size from 512 to 4096 while maintaining the flow batch size at 512, performing 8 parameter updates per OT batch. We continue using exact OT. Results appear in Tables 16 and 17.
Table 16: Generative Perplexity: Exact OT (MMF) — 4096 OT, 512 dff
@@ D.2 Continuous-Time Markov Chains Over Finite-State Spaces (Discrete Diffusion)
For the choice $Q_{t}=\sigma^{{}^{\prime}}(t)Q$, where $\sigma$ is monotonically increasing, $\sigma(0)=0$ and $\lim_{t\rightarrow 1}\sigma(t)=T$, we have $s_{t}=e^{\sigma(t)Q}s_{0}$. Matrices $Q$ must satisfy the properties of transition-rate matrices (Suhov and Kelbert, 2008), that is, they have non-negative non-diagonal entries, and the elements in each column add to 0. The choice for $Q$ is made such that: $\boldsymbol{s}_{1}$ is an easy reference distribution to sample from and the matrix exponential $e^{\sigma(t)Q}$ is easy to calculate (Austin et al., 2021; Campbell et al., 2022). Unfortunately, these conditions greatly restrict the design space in this framework.
@@ 31 James Wood, 1925-2002
Listing 6: Generated text from the DFM-MMF model with multi-mask source distribution and normal training, with sequence length L=128.
@@
Listing 7: Generated text from the DFM-MMF-OT model with multi-mask source distribution and minibatch-OT training, with sequence length L=128.